# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ XV.3 Section - Summary
We have presented a new formulation of perturbation theory based upon a mean-field- Hamiltonian-approximation and have demonstrated that it is Borel-summable for arbitrary coupling-strength in the case of QAHO, SAHO and the QDWO, by obtaining very accurate results for the energy-spectrum. By this demonstration, the distinction between the‘perturbative’ and the ‘non-perturbative’-regimes appears to be an artifact based on SFPT.
The results obtained for the QDWO may be of particular significance because this case typically belongs to systems with ‘degenerate ground states’, which are generically found to frustrate Borel-summability in the SFPT. It may also be significant that, results of comparable accuracy have also been obtained [104]( see, Section-VIII ) with a much cruder approximation using the square-well-potential in the ’mean-field Hamiltonian MFH’ which preserves the general features of the ’mean field approximation scheme (MFAS) ’, thus demonstrating the general nature and flexibility of the input in MFAS.
In view of the simplicity and the general-nature of the formulation demonstrated in this investigation , we are inclined to conjecture that MFPT may be applicable to arbitrary interacting systems in quantum-theory.
@@ XVI Summary and Conclusions
In summary, a non-perturbative scheme of approximation in quantum theory, designated herein as: “ Non-perturbative General Approximation Scheme” (NGAS) is presented in this report. This scheme is in principle, applicable to general interacting systems described by a Hamiltonian and has the distinction of being simple in formulation, non-perturbative, self-consistent, flexible to incorporate different choice of approximating input and more importantly, admits of systematic improvement via a perturbation theory, which is valid for arbitrary strength of interaction.
The crucial feature of this scheme consists of finding a “mapping” which maps the “interacting system” on to an “exactly solvable” input-model, while preserving the major effects of interaction through a self consistent feed-back mechanism of demanding equal quantum averages of observables in the two systems.
This approximation method has the advantage over the standard formulation of perturbation theory (SFPT) and the variational approximation by transcending the limitations of both: unlike the variational method, it is systematically improvable through the development of a new perturbation theory namely, the “Mean Field Perturbation Theory”(MFPT), whereas in contrast to the case of the SFPT, MFPT preserves the analyticity and non-linearity of the original system order-by-order, while being Borel-summable, for all allowed values of the coupling strength of interaction. Moreover, better behavior of the divergent asymptotic perturbation series at large orders, is seen to emerge in MFPT.
The scheme reproduces the results obtained by several earlier methods, e.g.[112-114] while transcending the limitations of these methods in respect of wider applicability, systematic improvement and better convergence.
@@
It may also be significant that NGAS allows flexibility in the choice of the approximating Hamiltonian (AH). Although the results presented here are based upon the “harmonic-approximation” , the same with comparable accuracy have also been obtained via a much cruder approximation using the square-well-Hamiltonian as the input Hamiltonian within the framework of the general features of NGAS.
@@
Similarly, in the domain of quantum field theory, the scheme has been applied to the case of $g\phi^{4}$ theory, with the choice of the approximating Hamiltonian as that for the free-hermitian-scalar field (but with adjustable ‘mass’ and ‘shifted’ field). In this case, the standard results of the Gaussian-approximation are reproduced in the LO including the non-perturbative renormalization of the ‘mass’ and ‘coupling strength’. However, NGAS appears to be more general in scope and insight than the Gaussian approximation since the former provides a ’dynamical explanation’ of the latter as emergent through the mechanism of altered vacuum-structure introduced by the interaction. Besides, the scheme goes beyond the scope of the Gaussian-approach in establishing new results, e.g., the calculation of the momentum-distribution of the condensate-structure function ‘$n(k)$’. It may also be emphasized that by going beyond the LO of NGAS, the results of the Gaussian approximation can be systematically improved, order-by-order.
The resulting momentum distribution of the vacuum condensate structure function ‘$n(k)$’ deserves special mention as it displays the non-standard feature of an appreciable spread in ${}|k|~{}$ about the origin, scaled by the renormalized mass of the physical quanta. It is reasonable to expect that this condensate-structure of the physical vacuum persists to exist at finite temperature manifesting in the thermodynamic properties of the system with testable consequences.
As has been noted above, a remarkable feature of NGAS is that, while the LO-results are essentially non-perturbative, yet the same are improvable order-by-order in MFPT. The other features of MFPT, which are distinct from those of the conventional perturbation method,e.g. the former is free from power-series-expansion in any physical parameter, including the coupling strength and applicable for arbitrary strength of interaction.Besides, it has been demonstrated that MFPT is Borel-summable (for arbitrary coupling-strength) in the case of QAHO, SAHO and the QDWO, by obtaining very accurate results for the energy-spectrum.
By circumventing the limitations of SFPT in MFPT, it is demonstrated that the distinction between the‘perturbative’ and the ‘non-perturbative’-regimes vanishes and, therefore the same could be a mere artifact based inthe former formulation i.e. the standard formulation of the perturbation theory (SFPT).
The demonstration of Bore-summability of MFPT for the QDWO may be of special significance and contrasted with those derived in SFPT, because this case typically belongs to systems with ‘degenerate ground states’ and such cases are generically found to frustrate Borel-summability in the SFPT. As discussed in this report, this situation , which has given rise to considerable efforts in recent literature [95-99] to surmount this problem of the SFPT for the QDWO and other such systems , is not warranted in MFPT.
In view of the simplicity and the general-nature of the formulation , it may be reasonable to conjecture that this scheme, NGAS may be applicable to arbitrary interacting systems in quantum-theory.
@@ Acknowledgments
The authors acknowledge useful comments and suggestions from the anonymous referees of the journals at references [48],[50],[82] and[110] . The computation and library facilities at the Institute of Physics, Bhubaneswar (IOPB) and NISER, Bhubaneswar are gratefully acknowledged where one of the authors , (BPM) held positions of guest-scientist and visiting faculty respectively. The authors are thankful to Prof. F.M.Fernandez for help in symbolic-computation using Maple and for helpful comments. Useful comments from Profs. A. M. Srivastava, S. G. Mishra and P. Agrawall during seminar talks on the topic(s) of this investigation are gratefully acknowledged.

### v3

@@ XV.3 Section - Summary
We have introduced a novel formulation of perturbation theory predicated on a mean-field Hamiltonian approximation, demonstrating its Borel-summability for arbitrary coupling strength in the contexts of QAHO, SAHO, and the QDWO, by achieving highly precise results for the energy spectrum. This demonstration suggests that the distinction between perturbative and non-perturbative regimes may be an artifact of SFPT.
The results obtained for the QDWO are particularly noteworthy, as this system typically features degenerate ground states, which generally impede Borel-summability in SFPT.
Given the simplicity and generality of the formulation demonstrated in this study, we are inclined to conjecture that MFPT may be applicable to arbitrary interacting systems in quantum theory.
@@ XVI Summary and Conclusions
In this report, we introduce a non-perturbative approximation scheme in quantum theory, termed the ”Non-perturbative General Approximation Scheme” (NGAS). This scheme is fundamentally applicable to general interacting systems described by a Hamiltonian and is characterized by its simplicity in formulation, non-perturbative nature, self-consistency, and flexibility to incorporate various choices of approximating inputs.
A key feature of this scheme is the identification of a ”mapping” that transforms the ”interacting system” into an ”exactly solvable” input model, while maintaining the primary effects of interaction through a self-consistent feedback mechanism that demands equal quantum averages of observables in both systems. This approximation method surpasses the standard formulation of perturbation theory (SFPT) and the variational approximation by overcoming their limitations: unlike the variational method, it is systematically improvable through the development of a new perturbation theory, namely, the ”Mean Field Perturbation Theory” (MFPT). Notably, it allows for systematic improvement through a perturbation theory that is valid for arbitrary interaction strengths. Besides, in contrast to SFPT, MFPT preserves the analyticity and non-linearity of the original system order-by-order, while being Borel-summable for all permissible values of the coupling strength of interaction. Furthermore, MFPT demonstrates improved behavior of the divergent asymptotic perturbation series at large orders. The scheme replicates results obtained by several earlier methods, e.g., [112-114], while surpassing these methods in terms of wider applicability, systematic improvement, and better convergence. The other features of MFPT, which distinguish it from the conventional perturbation methods, include its freedom from power-series expansion in any small parameter, including the coupling strength, and its applicability to arbitrary interaction strengths. Hence, by circumventing the limitations of SFPT in MFPT, it is established that the distinction between the ’perturbative’ and ’non-perturbative’ regimes vanishes, which suggests that this distinction may be a mere artifact of the former formulation, i.e., the standard formulation of perturbation theory (SFPT).
@@
It is noteworthy that NGAS allows flexibility in the choice of the approximating Hamiltonian (AH). Although the results presented here are based on the ”harmonic-approximation,” we have also demonstrated that comparable accuracy can be achieved using a much cruder approximation with the infinite- square-well-Hamiltonian as the input Hamiltonian within the framework of NGAS’s general features.
@@
In the realm of quantum field theory, the scheme has been applied to the $g\phi^{4}$ theory, with the approximating Hamiltonian selected as that for the free-hermitian-scalar field, albeit with adjustable mass and shifted field. In this context, the standard results of the Gaussian approximation are replicated at the leading order, including the non-perturbative renormalization of the mass and coupling strength. However, NGAS appears to offer a broader scope and deeper insight than the Gaussian approximation, as it provides a ’dynamical explanation’ of the latter, emerging through the mechanism of altered vacuum structure introduced by the interaction. Furthermore, the scheme extends beyond the Gaussian approach by establishing new results, such as the calculation of the momentum distribution of the condensate structure function ‘$n(k)$’. It is noteworthy that by advancing beyond the leading order of NGAS, the results of the Gaussian approximation can be systematically refined, order-by-order. The resulting momentum distribution of the vacuum condensate structure function ‘$n(k)$’ is particularly significant, as it exhibits the non-standard feature of a considerable spread in $|k|~$ about the origin, scaled by the renormalized mass of the physical quanta. It is plausible to anticipate that this condensate structure of the physical vacuum persists at finite temperature, manifesting in the thermodynamic properties of the system with testable consequences.
@@ Acknowledgments
Collaboration with Dr. Nabghan Santi and Dr. Noubihary Pradhanon on several aspects of the reasearch presented herein is acknowledged. The author acknowledges useful comments and suggestions from the anonymous referees of the journals at references [48],[50],[82] and[110] . The computation and library facilities at the Institute of Physics, Bhubaneswar (IOPB) and NISER, India are gratefully acknowledged where one of the authors , (BPM) held positions of guest-scientist and visiting faculty respectively. The author is thankful to Prof. F.M.Fernandez for help in symbolic-computation using Maple and for helpful comments. Useful remarks from Profs. A. M. Srivastava, S. G. Mishra and P. Agrawall during seminar talks on the topic(s) of this investigation are gratefully acknowledged.